# A Less Terrifying Universe? Mundanity as an Explanation for the Fermi Paradox

### v1

@@ 3 Colonizing the Galaxy
Thus, even with current technology, travel to the nearest stars may well be possible, even if exceptionally challenging and extremely slow. However, even if it is technically feasible to colonize the entire Galaxy, at least using self-replicating probes such as proposed by Hart (1975), for example, the colonizers will require a motivation to do this. The benefits obtained must outweigh the cost and potential risks. Since it is unlikely that there would be much benefit from bringing physical objects back, the benefits must come from the scientific knowledge acquired and then sent back to the home planet(s).
@@ 4 A Mundane Perspective
There are some indications that technological and engineering changes may already be becoming more modest. Park et al. (2023) argued that papers and patents are becoming less “disruptive” over time. Some have noted a slowing down of Moore’s law (e.g. Lundstrom & Alam, 2022; Theis & Wong, 2017). Ramifications may be seen in the flattening of the growth rate of “Top 500” supercomputer list. [^5] https://top500.org/statistics/perfdevel/ Even if this is just a temporary slowdown in technological progress, a fundamental limit would eventually be reached. Also, as an additional constraint, Haqq-Misra & Baum (2009) and Mullan & Haqq-Misra (2019) argue that sustainability places limits on the growth of civilizations, and Haqq-Misra et al. (2025) propose that stellar luminosity and mass provide a limit in the absence of very advanced technology.
@@ 4.2.1 Limits to Galactic Colonization, Even by Robots
For significant colonization or exploration of the Galaxy by sending probes to almost all star systems would again have to have benefits that outweigh the costs. Or, in terms of Keynes’ “animal spirits”, greed must overcome fear (Keynes 1936). Robotic exploration should be far cheaper, particularly if self-replicating von Neumann probes (Freitas, 1980) are sent, where only a few initial probes need to be sent out. The benefits depend on whether much is learned from each in situ exploration of a star system. Probably not much new scientific information would be provided by directly visiting, for example, many thousands of M stars, let alone billions. The cost would then be that of the initial launch, which potentially is not too great amortized over the entire lifetime of the civilization. However, there may be dangers in creating a vast number of (probably) intelligent devices. This is the “Berserker” hypothesis of deadly probes as described in the science fiction novels of Fred Saberhagen and discussed by, for example, Brin (1983) and citetCirkovic2009. While the concept of Berserker probes may be extreme, the fear of this happening can be regarded as unexceptional. Indeed, this is something that concerns contemporary society with the development of AI - see for example Bostrom (2014), Amodei et al. (2016), Joy (2020), and Müller & Bostrom (2025). In the mundane perspective, where other civilizations, are not that much more advanced, a limit to exploration would arise when it was found that not much new was found from each encounter or other costs are incurred. In this case the benefit to risk ratio would decrease with each encounter and the passage of time.
@@ 6.3 Nearby Artifacts - the Solar System, and UFOs/UAPs on the Earth
In our Solar System, perhaps the most dramatic, and controversial, claim for the detection of an alien artifact is the first interstellar object to be identified passing through the Solar System - ‘Oumuamua (1I/2017 U1). Loeb (2022) has proposed that ‘Oumuamua may have an artificial origin, perhaps as a fragment of a disintegrating Dyson sphere (Loeb, 2023). This has been disputed by, for example Zhou et al. (2022) and Zuckerman (2022). Further, Davenport et al. (2025) caution that care must be taken when comparing interstellar bodies with solar system objects.
@@ 8 What if One or Both Mundane Premises are False?
Very Few Civilizations, Very High Technology Level: A solitary civilization would have a strong motivation to detect other ETCs. With ETCs being rare, each new one would be of great interest. Possessing very advanced technology would enable it to create high-powered beacons, and spread robotic probes throughout the Galaxy. We then run into the Fermi paradox given our lack of the detection of probes. Also, this would be inconsistent with the Copernican mediocrity principle as we would be one of a very few Technological Civilizations.
@@
Very Few Civilizations, Mundane Technology Level: This situation naturally explains the Great Silence. However, it also invalidates one of the assumption behind the Fermi paradox, that a reasonable number of ETCs would arise and rather rapidly spread throughout the Galaxy. Although it might be viewed as an extreme version of the mundanity paradigm, there would still be a drive for such civilizations to explore large volumes of the Galaxy. As noted in Section 3, this is perhaps achievable by near-future technology on the Earth.
@@ 9 Conclusion
I thank Jacob Haqq-Misra and Ravi Kumar Kopparapu and for useful comments. The work was supported in part by NASA under award number 80GSFC21M0002.

### v2

@@ 3 Colonizing the Galaxy
Thus, even with current technology, travel to the nearest stars may well be possible, even if exceptionally challenging and extremely slow. However, even if it is technically feasible to colonize the entire Galaxy, at least using self-replicating probes such as proposed by Hart (1975), for example, the colonizers will require a motivation to do this. The benefits obtained must outweigh the costs and potential risks. Since it is unlikely that there would be much benefit from bringing physical objects back, the benefits must come from the scientific knowledge acquired and then sent back to the home planet(s).
@@ 4 A Mundane Perspective
There are some indications that technological and engineering changes may already be becoming more modest. Park et al. (2023) argued that papers and patents are becoming less “disruptive” over time. Some have noted a slowing down of Moore’s law (e.g. Lundstrom & Alam, 2022; Theis & Wong, 2017). Ramifications may be seen in the flattening of the growth rate of “Top 500” supercomputer list. [^5] https://top500.org/statistics/perfdevel/ Even if this is just a temporary slowdown in technological progress, a fundamental limit would eventually be reached. Also, as possible additional constraints, Haqq-Misra & Baum (2009) and Mullan & Haqq-Misra (2019) argue that sustainability places limits on the growth of civilizations, Haqq-Misra et al. (2025) propose that stellar luminosity and mass provide a limit in the absence of very advanced technology, and Balbi & Lingam (2025) consider constraints imposed by production of waste heat.
@@ 4.2.1 Limits to Galactic Colonization, Even by Robots
For significant colonization or exploration of the Galaxy by sending probes to almost all star systems would again have to have benefits that outweigh the costs. Or, in terms of Keynes’ “animal spirits”, greed must overcome fear (Keynes 1936). Robotic exploration should be far cheaper, particularly if self-replicating von Neumann probes (Freitas, 1980) are sent, where only a few initial probes need to be sent out. The benefits depend on whether much is learned from each in situ exploration of a star system. Probably not much new scientific information would be provided by directly visiting, for example, many thousands of M stars, let alone billions. The cost would then be that of the initial launch, which potentially is not too great amortized over the entire lifetime of the civilization. However, there may be dangers in creating a vast number of (probably) intelligent devices. This is the “Berserker” hypothesis of deadly probes as described in the science fiction novels of Fred Saberhagen and discussed by, for example, Brin (1983) and Cirkovic (2009). While the concept of Berserker probes may be extreme, the fear of this happening can be regarded as unexceptional. Indeed, this is something that concerns contemporary society with the development of AI - see for example Bostrom (2014), Amodei et al. (2016), Joy (2020), and Müller & Bostrom (2025). In the mundane perspective, where other civilizations are not that much more advanced, a limit to exploration would arise when it was found that not much new was found from each encounter or other costs are incurred. In this case the benefit to risk ratio would decrease with each encounter and the passage of time.
@@ 6.3 Nearby Artifacts - the Solar System, and UFOs/UAPs on the Earth
In our Solar System, perhaps the most dramatic, and controversial, claim for the detection of an alien artifact is the first interstellar object to be identified passing through the Solar System - ‘Oumuamua (1I/2017 U1). Loeb (2022) has proposed that ‘Oumuamua may have an artificial origin, perhaps as a fragment of a disintegrating Dyson sphere (Loeb, 2023). This has been disputed by, for example, Zhou et al. (2022) and Zuckerman (2022). Further, Davenport et al. (2025) caution that care must be taken when comparing interstellar bodies with solar system objects.
@@ 8 What if One or Both Mundane Premises are False?
Very Few Civilizations, Very High Technology Level: A solitary civilization would have a strong motivation to detect other ETCs. With ETCs being rare, each new one would be of great interest. Possessing very advanced technology would enable it to create high-powered beacons, and spread robotic probes throughout the Galaxy. We then run into the Fermi paradox given our lack of the detection of probes. In this case there would finally be a breakdown of the Copernican mediocrity principle, as we would be one of a very small number of Technological Civilizations.
@@
Very Few Civilizations, Mundane Technology Level: This situation naturally explains the Great Silence. However, it also invalidates one of the assumption behind the Fermi paradox, that a reasonable number of ETCs would arise and rather rapidly999i.e., on a timescale much less than the age of the Milky Way spread throughout the Galaxy. Although it might be viewed as an extreme version of the mundanity paradigm, there would still be a drive for such civilizations to explore large volumes of the Galaxy. As noted in Section 3, this is perhaps achievable by near-future technology on the Earth.
@@ 9 Conclusion
acknowledgements
I thank Jacob Haqq-Misra, Ravi Kumar Kopparapu, Viktor Sandner and an anonymous referee for useful comments. The work was supported in part by NASA under award number 80GSFC21M0002.